OpenStudy (anonymous):

(x + 3)^1/2 - 1 = x All help is appreciated! :)

OpenStudy (anonymous):

what about it

OpenStudy (anonymous):

Find x.

OpenStudy (anonymous):

Check for extraneous solutions.

OpenStudy (anonymous):

-1 part of the exponent?

OpenStudy (anonymous):

Nope! Sorry..didn't clarify.

OpenStudy (anonymous):

@petewe

OpenStudy (anonymous):

move 1 over, square the equation, solve for x

OpenStudy (lauluisa):

It's equal to 1.

OpenStudy (anonymous):

Thanks! How did you find it?

OpenStudy (lauluisa):

(x+3)^1 It-s always going to be the same number because u r multiplying. So... I skiped that step. which was easier. x+3/2-1=x x/2 + 3/2 - 1 = X Now u substract x/2 from both sides and u r going to end up with: 1/2= x-x/2 1/2 = 1/2x Multiply 2 in both sides.. x=1

OpenStudy (anonymous):

Thank you so much!

OpenStudy (lauluisa):

Anytime :)

OpenStudy (anonymous):

Question, when you solved the (x+3)^1 part, did you know that the ^1 was supposed to be ^1/2? Here it is in a pic. I am just double checking cuz you just said (x+3)^1, so I wasn't sure. Just making sure! Thanks a bunch. @lauluisa